# CaVE: A Cone-Aligned Approach for Fast Predict-then-optimize with Binary Linear Programs

### 12 Dec 2023 v1

@@ 7 Conclusion
The best of the three seems to be CaVE+, which regresses on an inner vector of the optimal (sub)cone of a training instance, resulting in stable and efficient training, as well as test regret results that compare to state-of-the-art methods that require training 30 times longer on CVRP20, or do not even complete a single training epoch in 10 hours for CVRP30. We note that if we were to use early termination during training, CaVE methods would record even smaller training times as they do converge in fewer gradient descent iterations than competing methods as per Fig. 4. We hope that our framework will enable the adoption of end-to-end predict-then-optimize in a wider range of applications and have made our implementation available with plans to make CaVE one of the standard methods within the PyEPO package following the publication of this work.
CaVE is limited to binary problems. In practice, this is not too problematic, as bounded integer variables can be represented using a set of binary variables. Another limitation of our method is the lack of theoretical guarantees. In particular, we currently do not know whether the loss function  CaVE or a modification thereof could be proven to be a valid upper bound on regret, as does the SPO+ loss of Elmachtoub and Grigas [7] for example. This direction merits further investigation. An interesting connection may be established between CaVE and recent ML methods such as [19] for predicting the active constraints of a family of similar optimization problems for which optima are known. Rather than predict-then-optimize, the goal here is to predict the active set accurately to solve a reduced optimization problem over only that set.
@@ Appendix 0.A More Experiments
For enhancing training efficiency, it has been proposed by [17] that linear relaxation can be used as a substitute for the original BLP during training. We applied linear relaxation to the Miller-Tucker-Zemlin (MTZ) formulation [18], and trained SPO+ Rel and PFYL Rel on the same TSP instances with 20 and 50 nodes. Although relaxation methods are more efficient than CaVE on TSP20, they yield higher regret. As the model size increases, relaxation approach also lose their efficiency advantage.
@@
(a) Average Test Normalized Regret
(b) Average Training Time (Sec)
@@
Deg $4$ 50.38 54.91
Deg $6$ 73.79 69.75
(b) Average Training Time (Sec)
@@
(a) Average Test Normalized Regret
(b) Average Training Time (Sec)
@@
Deg $4$ 530.63 521.95
Deg $6$ 528.50 529.40
(b) Average Training Time (Sec)
AGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 15 Mar 2024 v2

@@ 7 Conclusion
The best of the three appears to be CaVE+, which regresses on an inner vector of the optimal subcone of a training instance, resulting in stable and efficient training, as well as test regret results that compare to state-of-the-art methods that require training 30 times longer on CVRP20, or do not even complete a single training epoch in 8 hours for CVRP30. We note that if we were to use early termination during training, CaVE methods would record even smaller training times as they do converge in fewer gradient descent iterations than competing methods as per Fig. 4. We hope that our framework will enable the adoption of end-to-end predict-then-optimize in a wider range of applications and have made our implementation available with plans to make CaVE one of the standard methods within the PyEPO package following the publication of this work.
CaVE is limited to binary problems. In practice, this is not too problematic, as bounded integer variables can be represented using a set of binary variables. Another limitation of our method is the lack of theoretical guarantees. In particular, we currently do not know whether the loss function in Eq. (4) or a modification thereof could be proven to be a valid upper bound on regret, as does the SPO+ loss of Elmachtoub and Grigas [7] for example. This direction merits further investigation. An interesting connection may be established between CaVE and recent ML methods such as [19] to predict the active constraints of a family of similar optimization problems for which optima are known. Rather than predict-then-optimize, the goal of Misra et al. [19] is to accurately predict the active set to solve a reduced optimization problem over only that set.
@@ Appendix 0.A More Experiments
To enhance training efficiency, it has been proposed in [17] that a linear relaxation can be used as a substitute for the original BLP during training. Because the DFJ formulation utilizes constraint generation to handle subtour elimiation constraints, it is challenging to achieve linear relaxation due to the exponential number of constraints. In this study, we used the LP relaxation of the Miller-Tucker-Zemlin (MTZ) formulation [18] of the TSP, and trained SPO+ Rel and PFYL Rel on the same TSP instances with 20 and 50 nodes. Although relaxation methods are more efficient than CaVE in TSP20, they result in higher regret. As the size of the model increases, the relaxation approach also loses its efficiency advantage.
@@
(a) Average Test Normalized Regret (%) with Standard Deviation
(b) Average Training Time (Sec) with Standard Deviation
@@
Deg $4$ $71.92\pm 2.05$ $77.25\pm 0.60$
Deg $6$ $67.99\pm 0.60$ $53.32\pm 3.74$
(b) Average Training Time (Sec) with Standard Deviation
@@
(a) Average Test Normalized Regret (%) with Standard Deviation
(b) Average Training Time (Sec) with Standard Deviation
@@
Deg $4$ $386.06\pm 9.69$ $536.67\pm 4.94$
Deg $6$ $636.99\pm 3.04$ $510.37\pm 3.46$
(b) Average Training Time (Sec) with Standard Deviation
0.A.2 CVRP Relaxation.
Similarly, the CVRP formulation with k-path cuts also struggles to obtain a linear relaxation. Thus, the MTZ formulation can be modified with constraints
$u_{j}-u_{i}\geq Q(x_{ij}-1)+q_{j}\quad\forall i\neq j,i\neq 0,j\neq 0$
for subtour elimilation and customer demand, where $Q$ is the capacity and $q_{i}$ is the demand of customer $i$. The SPO+ Rel and PFYL Rel are trained for the same VRP instances with 20 customers. While the linear relaxation approach demonstrates high efficiency in these scenarios, it does not guarantee a low regret compared to CaVE.
Table 9: Experimental Results for CVRP20202020 Relaxation
(a) Average Test Normalized Regret (%) with Standard Deviation
(b) Average Training Time (Sec) with Standard Deviation
Methods SPO+ Rel PFYL Rel
Deg $4$ $8.03\pm 0.38$ $17.07\pm 0.63$
Deg $6$ $15.73\pm 0.39$ $19.19\pm 1.66$
Methods SPO+ Rel PFYL Rel
Deg $4$ $78.95\pm 0.73$ $78.80\pm 1.19$
Deg $6$ $78.74\pm 3.82$ $81.80\pm 0.86$
(b) Average Training Time (Sec) with Standard Deviation